OpenStudy (anonymous):

still cant get the correct answer. Minimize F=x^2+y^2 with x + 2y = 10.

OpenStudy (agent0smith):

I think you want to use substitution - using x+2y =10, so x = 10-2y replace an x in the F equation with that, so \[\large F = (10-2y)^2+y^2\] Try expanding and simplifying.

OpenStudy (anonymous):

still got the wrong answer

OpenStudy (agent0smith):

Don't just tell us you got the wrong answer... show your work so we can see where you're going wrong.